Calculate the electric flux through a rectangular surface of dimensions 20cm × 10cm, placed with its positive normal making an angle of 30° with a uniform electric field of 30 × 10^-3 N/C.

1
Expert's answer
2022-03-09T10:33:06-0500

The electric flux

"\Phi=EA\cos\theta\\
=30*10^{-3}\:{\rm N/C*0.2*0.1\: m^2}*\cos 30^\circ\\
=5.2*10^{-4}\:\rm N\cdot m^2/C"